### v2

# A note on the planar triangles in Minkowski spacetime

Abstract
The geometry of 2D Minkowski spacetime $\mathbb{R}^{1,1}$ (or Minkowski plane [^1] This is not to be confused with the 2D non-Euclidean normed space [5], also called the Minkowski plane.) is similar but fundamentally different from the more familiar Euclidean plane geometry [1, 2, 3, 4]. This note gives an elementary discussion on some basic properties of a triangle on the Minkowski plane. In particular, we show that the theorem of Feuerbach [^2] In 1822, Feuerbach discovered the nine-point circle and also proved that it is internally tangent to the incircle and externally tangent to the excircles of the triangle [6], see [7, 8, 9, 5, 10, 11] for some extensions of this theorem within or beyond the Euclidean plane geometry. also holds and a use of the incenter/excenters is pointed out.
Taking the speed of light $c=1$, the Minkowski plane can be described by the line element $ds^{2}=dx^{2}-dt^{2}$ with a set of spacetime coordinate $(x,t)$, it is both flat and translationally invariant, thus has straight geodesics; physically these correspond to the worldlines of constant-velocity free-falling observers, which can be timelike ($|\Delta t/\Delta x|>1$), null ($|\Delta t/\Delta x|=1$) or spacelike ($|\Delta t/\Delta x|<1$). We define the distance between two points $(x_{1},t_{1})$ and $(x_{2},t_{2})$ (or the length of a vector $(x_{2}-x_{1},t_{2}-t_{1})$ connecting them) as $\sqrt{(x_{1}-x_{2})^{2}-(t_{1}-t_{2})^{2}}$, with positive real or imaginary part. One can also use a polar chart $(r,\theta)$ defined by $\vec{r}=(x,t)=r(\sinh\theta,\cosh\theta)$ with $r,\theta\in\mathbb{R}$ if $\vec{r}$ is timelike, the length $|\vec{r}|$ of the vector $\vec{r}$ is then $i|r|$. Given two vectors $\vec{a}=(x_{a},t_{a})$ and $\vec{b}=(x_{b},t_{b})$, a natural definition for their inner product (from the metric) is $\vec{a}\cdot\vec{b}=x_{a}x_{b}-t_{a}t_{b}$, such that $|\vec{r}|^{2}=\vec{r}\cdot\vec{r}$. Using this inner product, the angle between two timelike vectors can be defined as $\angle(\vec{a},\vec{b})=-\text{sgn}(\vec{a}\cdot\vec{b})\,\text{arccosh}\left[\vec{a}\cdot\vec{b}/(|\vec{a}|\,|\vec{b}|)\right]$ (with $\text{arccosh}\,x>0$), such that for $\vec{a}=r_{a}(\sinh\theta_{a},\cosh\theta_{a})$ and $\vec{b}=r_{b}(\sinh\theta_{b},\cosh\theta_{b})$, $\angle(\vec{a},\vec{b})=|\theta_{a}-\theta_{b}|\,\text{sgn}(r_{a}r_{b})$, since $\vec{a}\cdot\vec{b}<0$. The angle between a timelike vector $\vec{a}$ and a spacelike vector $\vec{b}$ can be defined as $\angle(\vec{a},\vec{b})=\text{arcsinh}\left[i(\vec{a}\cdot\vec{b})/(|\vec{a}|\,|\vec{b}|)\right]$, such that for $\vec{a}=r_{a}(\sinh\theta_{a},\cosh\theta_{a})$ and $\vec{b}=r_{b}(\cosh\theta_{b},\sinh\theta_{b})$, $\angle(\vec{a},\vec{b})=(\theta_{a}-\theta_{b})\,\text{sgn}(r_{a}r_{b})$. Both the distance and the included angle are invariant under the coordinate transformation so called Lorentz boost, e.g.,
$\left(\begin{matrix}t^{\prime}\\ x^{\prime}\end{matrix}\right)=e^{\alpha\left(\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}\right)}\left(\begin{matrix}t\\ x\end{matrix}\right)=\frac{\left(\begin{smallmatrix}1&\beta\\ \beta&1\end{smallmatrix}\right)}{\sqrt{1-\beta^{2}}}\left(\begin{matrix}t\\ x\end{matrix}\right)=\left(\begin{matrix}\cosh\alpha&\sinh\alpha\\ \sinh\alpha&\cosh\alpha\end{matrix}\right)\left(\begin{matrix}r\cosh\theta\\ r\sinh\theta\end{matrix}\right)=\left(\begin{matrix}r\cosh\theta^{\prime}\\ r\sinh\theta^{\prime}\end{matrix}\right),$ (1)
(where $\beta$ is the boost’s velocity and $\alpha$ is the so called rapidity) which leads to a hyperbolic rotation: $\theta^{\prime}=\theta+\alpha$, while preserving the inner product.
Two vectors are orthogonal to each other if their inner product vanishes; a timelike vector is orthogonal to a spacelike vector, while a null vector is orthogonal to itself. Given three points $\vec{A}=(x_{A},t_{A})$, $\vec{B}=(x_{B},t_{B})$, $\vec{C}=(x_{C},t_{C})$, it is straightforward to show that there is a unique point $H_{A}$ on the line $BC$ satisfying $(\vec{H}_{A}-\vec{A})\cdot(\vec{B}-\vec{C})=0$, which is the perpendicular feet dropped from $A$ on $BC$. The distance between $A$ and the line $BC$, defined as the length of $\vec{H}_{A}-\vec{A}$, is given explicitly by
$|\vec{H}_{A}-\vec{A}|=\sqrt{-\frac{[x_{B}(t_{C}-t_{A})+x_{C}(t_{B}-t_{A})+x_{A}(t_{B}-t_{C})]^{2}}{(x_{B}-x_{C})^{2}-(t_{B}-t_{C})^{2}}}.$ (2)
If the distances between a point $C^{\prime}$ and both $CB$ and $CA$ are same, then $C^{\prime}$ lies in the bisector of $\angle(\vec{A}-\vec{C},\vec{B}-\vec{C})$, this bisector can be constructed as follows. First we find a point $\vec{r}_{*}=(x_{*},t_{*})$ on the line $BC$ satisfying $|\vec{r}_{*}-\vec{C}|=|\vec{A}-\vec{C}|$, the line passing through both $\vec{C}$ and $(\vec{A}+\vec{r}_{*})/2$ then gives the bisector, explicitly we have
$x_{*}^{\pm}=x_{c}\pm(x_{B}-x_{C})\sqrt{\frac{|\vec{A}-\vec{C}|^{2}}{|\vec{B}-\vec{C}|^{2}}},\quad t_{*}^{\pm}=\frac{t_{B}-t_{C}}{x_{B}-x_{C}}(x_{*}^{\pm}-x_{C})+t_{C}.$ (3)
The bisector of $\angle(\vec{A}-\vec{C},\vec{B}-\vec{C})$ is $\frac{t-t_{C}}{x-x_{C}}=\frac{(t_{*}^{+}+t_{A})/2-t_{C}}{(x_{*}^{+}+x_{A})/2-x_{C}}\equiv f_{+}(A,B,C)$, while the bisector of its supplementary angle is $\frac{t-t_{C}}{x-x_{C}}=\frac{(t_{*}^{-}+t_{A})/2-t_{C}}{(x_{*}^{-}+x_{A})/2-x_{C}}\equiv f_{-}(A,B,C)$. The incenter of triangle $\triangle ABC$ is given by the intersection point of three inner bisectors, namely
$\frac{t-t_{C}}{x-x_{C}}=f_{+}(A,B,C),\quad\frac{t-t_{A}}{x-x_{A}}=f_{+}(B,C,A),\quad\frac{t-t_{B}}{x-x_{B}}=f_{+}(C,A,B).$ (4)
It is inside the triangle. As can be seen from Eq. (3), only when $|\vec{A}-\vec{C}|^{2},|\vec{A}-\vec{B}|^{2},|\vec{B}-\vec{C}|^{2}$ have the same sign does the incenter exist, that is, we can only talk about the incenter of a timelike/spacelike triangle, the all sides of which are timelike/spacelike.
The Minkowski plane can be mapped into a Euclidean plane by simply identifying $(x,t)$ with the Euclidean coordinate $(x,y)$, which is the Minkowski diagram. Under this mapping, a straight line remains straight, a Minkowski circle [^3] Physically, a Minkowski circle $x^{2}-t^{2}=r^{2}$ with real radius $r$ is associated with a Rindler observer with constant proper acceleration $a=1/r$, i.e., $\frac{d}{dt}\left(\frac{\dot{x}}{\sqrt{1-\dot{x}^{2}}}\right)=a$, the solution can be parameterized by the proper time $d\tau=\sqrt{1-\dot{x}^{2}}\,dt$ as $x(\tau)=x_{0}+a^{-1}[\cosh(a\tau)-1]$, $t(\tau)=t_{0}+a^{-1}\sinh(a\tau)$, and therefore $(x-x_{0}+a^{-1})^{2}-(t-t_{0})^{2}=a^{-2}$. (the curve with constant distance to a point), however, turns into a Euclidean equilateral hyperbola. Once we draw a Minkowski triangle $\triangle ABC$, just as the Euclidean case, there are several Minkowski circles associated with it, these include the incircle (with radius $r_{\star}$) centered at the incenter $I$, the excircles (with radius $r_{a,b,c}$) centered at the excenters $I_{a,b,c}$ opposite to the vertex $A,B,C$, respectively, and the circumcircle [^4] It is easy to show that the circumcircle exists for a generic triangle without null sides. (with radius $R$) centered at the circumcenter $O$, specifically we choose $\vec{O}=(0,0)$ such that $|\vec{A}|=|\vec{B}|=|\vec{C}|=R$. The definitions of these circles are akin to their Euclidean versions, only with the condition of tangency defined in the Minkowski sense, and the incenter/excenter exists only for a non-mixed (i.e., timelike or spacelike) triangle. We denote the sidelengths of the triangle by $a=|\vec{B}-\vec{C}|$, $b=|\vec{A}-\vec{C}|$, $c=|\vec{A}-\vec{B}|$, and the semiperimeter by $s=(a+b+c)/2$. Perhaps not surprisingly, Heron’s formula $\Delta=\sqrt{s(s-a)(s-b)(s-c)}$ is still valid, where $\Delta$ is the area of $\triangle ABC$ defined by $\Delta=\frac{1}{2}\times\text{Base}\times\text{Height}=\frac{1}{2}|\vec{B}-\vec{C}||\vec{H}_{A}-\vec{A}|=\frac{i}{2}|(x_{B}-x_{A})(t_{C}-t_{A})-(t_{B}-t_{A})(x_{C}-x_{A})|$. Moreover, one finds that
$\vec{I}=\frac{a}{2s}\vec{A}+\frac{b}{2s}\vec{B}+\frac{c}{2s}\vec{C},\quad\vec{I}_{c}=\frac{a}{2(s-c)}\vec{A}+\frac{b}{2(s-c)}\vec{B}-\frac{c}{2(s-c)}\vec{C},$ (5)
with
$r_{\star}=\frac{\Delta}{s},\quad r_{c}=\lambda_{c}\frac{\Delta}{s-c},\quad R=-\frac{abc}{\Delta},\quad\lambda_{c}=\text{sgn}(|a|+|b|-|c|).$ (6)
In comparison, the Euclidean result [12] is $r_{c}=\frac{\Delta}{s-c}$ with $R=\frac{abc}{\Delta}$. The area is imaginary, for a timelike triangle, the side lengths are imaginary hence $\{R,r_{\star},r_{c}\}$ are real. The orthocenter $H$ is given by $\vec{H}=\vec{A}+\vec{B}+\vec{C}$, and for a non-mixed triangle turns out to be the excenter of its pedal triangle opposite to $H_{\Lambda}$, where $\Lambda$ is the only side with $\lambda_{\Lambda}=-1$; for a mixed triangle, $\lambda$ can be positive for all three sides, in this case $\vec{H}$ is the incenter of its pedal triangle. Finally, the nine-point circle — a circle that passes through $H_{A,B,C}$,$\frac{\vec{A}+\vec{B}}{2}$, $\frac{\vec{A}+\vec{C}}{2}$, $\frac{\vec{B}+\vec{C}}{2}$, $\frac{\vec{A}+\vec{H}}{2}$, $\frac{\vec{B}+\vec{H}}{2}$, $\frac{\vec{C}+\vec{H}}{2}$ — exists, it is centered at $\vec{N}=\frac{1}{2}\vec{H}$ (on the Euler line) and has radius $R/2$. Following the same approach of [12], we can obtain
$|\vec{I}-\vec{N}|=R/2+r_{\star},\quad|\vec{I}_{c}-\vec{N}|=\sqrt{(R/2-\lambda_{c}r_{c})^{2}},$ (7)
which gives the Feuerbach theorem on Minkowski plane. In comparison, the Euclidean result is $|\vec{I}-\vec{N}|=R/2-r_{\star}$, $|\vec{I}_{c}-\vec{N}|=R/2+r_{c}$. Examples of Minkowski triangle and its nine-point circle are depicted in Fig. 1.
Figure 1: Minkowski triangle mapped into a Euclidean plane. Left: a timelike triangle. Right: a spacelike triangle. The red and blue dashed curves correspond to the incircle (centered at $I$) and an excircle (centered at $I_{b}$), respectively; the yellow lines are null trajectories passing through $I$; the gray curve is the nine-point circle (centered at $N$). Actually by switching the $x$ and $t$ coordinates we get a spacelike diagram from a timelike diagram, they are symmetrical about the line $x=t$. If $A,B,C$ are not at the same branch of the circumcircle (green curve), we get a mixed triangle.
The incenter and excenters of a non-mixed Minkowski triangle have a simple physical interpretation. The incircle is tangent to all three sides of the triangle (or their extensions), if we denote the points of tangency by $T_{A,B,C}$, then $|\vec{T}_{A}-\vec{I}|=|\vec{T}_{B}-\vec{I}|=|\vec{T}_{C}-\vec{I}|=r_{\star}$. Consider three ships with constant colinear velocity whose worldlines form a timelike triangle $\triangle ABC$, if a light source is placed on its incenter, the light pulse would take exactly same amount of time to reach these ships in their comoving frames, which is $\Delta t=r_{\star}$. One can also conceive the inverse process. The situation is similar for the three excenters.
Some of the results presented above for the triangles (2-simplices) can possibly be extended to the higher dimensional Minkowski spacetime $\mathbb{R}^{1,n}$, where the Minkowski sphere with imaginary radius is a hyperbolic space $\mathbb{H}^{n}$, and the sphere with real radius $1/H$ is a de Sitter spacetime $\text{dS}^{n}$.
Appendix A Pseudo-Complex Representation of $H_{A}$
Just as a point $(x,y)$ on the Euclidean plane can be represented by a complex number $z=x+iy$ such that $zz^{*}=x^{2}+y^{2}$, a point $(x,t)$ on the Minkowski plane can be represented by a pseudo-complex number (see for example [13, 3]) $z=x+\mathcal{I}t$ with $z^{*}=x-\mathcal{I}t$ and $\mathcal{I}^{2}=1$ such that $zz^{*}=x^{2}-t^{2}$ and $\cosh\theta+\mathcal{I}\sinh\theta=e^{\mathcal{I}\theta}$.
In this pseudo-complex (pc) representation, we can write $A=x_{A}+\mathcal{I}t_{A}$, $B=x_{B}+\mathcal{I}t_{B}$, $C=x_{C}+\mathcal{I}t_{C}$ with the circumcenter $O=0$, then for circumradius $R=\sqrt{\pm 1}$, $H_{A}=\frac{H\mp A^{*}BC}{2}$, where $H=A+B+C$. This can be proved by showing that $\text{Im}[(B-H_{A})(C-H_{A})^{*}]=\text{Re}[(B-H_{A})(A-H_{A})^{*}]=0$, where Re/Im denotes the real/imaginary part of a pc number.